OpenStudy (anonymous):

What is the exact value of the square root of 54 over 3 the square root 2.? A:3 the square root of 2 all over 2. B:9 the square root of 2 over 2. C:the square root of 3. D:33 the square root of 3.

OpenStudy (campbell_st):

\[\sqrt{54} = 3\sqrt{6}\] use this simplification with the rest of the question

OpenStudy (anonymous):

I thought it would be 3 square root 2

OpenStudy (campbell_st):

here is another clue \[\frac{3\sqrt{6}}{3\sqrt{2}}\]